OpenStudy (anonymous):

find the point on the line x + 2y = 2 that is closest to the origin, by the method of elimination and also by lagrange's method. Prove that your result is indeed a minimum.

OpenStudy (amistre64):

determine the perp line that goes through the origin; and see where they cross

OpenStudy (amistre64):

not quite sure what a lagrange method is :/

OpenStudy (amistre64):

x + 2y = 2 2x - y = 0 x + 2y = 2 4x - 2y = 0 ----------- 5x = 2 , x = 2/5, therefore y = 4/5

OpenStudy (anonymous):

@mahmit2012 dudeee ?

OpenStudy (anonymous):

@Hero dude ??

hero (hero):

What's wrong with what @amistre64 showed?

OpenStudy (anonymous):

@Hero it is absolutely not about lagrange method :/

hero (hero):

I'm not familiar with lagrange method either

OpenStudy (anonymous):

okey then ./

hero (hero):

If I had time, I could look into it for you, but I don't have time

OpenStudy (anonymous):

thanks again